OpenStudy (anonymous):

find the factored form of 7w-w^2=0 explain.

OpenStudy (anonymous):

u would google lol

OpenStudy (anonymous):

lol does google really have the answer?

OpenStudy (anonymous):

\[7w-w^2=0\\w(7-w)=0\\w=0 \vee w=7\]

OpenStudy (anonymous):

yeah google has it ..

OpenStudy (anonymous):

lol^

OpenStudy (anonymous):

im still lost cause the choices are w(7-w)=0 7(w-w^2)=0 w(7-w^2)=0

OpenStudy (anonymous):

it a .....

OpenStudy (anonymous):

hellpp, im a cute girl in need lol.

OpenStudy (anonymous):

i told u it was a .....u gonna message me back?